OpenStudy (anonymous):

If anyone could help I'd be really grateful √(3)*sinx+cosx=2

OpenStudy (anonymous):

ouch -.-

OpenStudy (anonymous):

help you do what?

OpenStudy (anonymous):

Solve the equation and decide what x should be element of.. sorry if poorly translated.

OpenStudy (anonymous):

i dont understand the question man

OpenStudy (raden):

a sin(x) + b cos(x) can be rewriten as k cos(x-θ) with k = sqrt(a^2 +b^2) and tanθ =a/b (we can solve for θ)

OpenStudy (amoodarya):

put in equation tan=tg pi/3 instead of sqrt3